OpenStudy (anonymous):

Jack usually mows his lawn in 4 hours. Marilyn can mow the same yard in 7 hours. How much time would it tak for them to mow the lawn together?

OpenStudy (anonymous):

Is the answer 3.23 hours

OpenStudy (ybarrap):

j = Jack mows 1 yard in 4 hours = 1/4 yards per hour m = Marilyn mows 1 yard in 7 hours = 1/6 yards per hour t = time for both to mow 1 yard So, (1/4) t + (1/7) t = 1 yard , t(1/4 + 1/7) = 1 t = 1/(1/4 + 1/7) = 1/11/28 = 28/11 = 2.54 hours Note. In 2.54 hours, jack cuts 2.54 * (1/4) = 0.635 of a yard while Marilyn cuts 2.54 * (1/7) = 0.363 of a yard. Together in 2.54 hours, they cut .635 + .363 = 1 yard. Ans. 2.54 hours